# Towards Definition of Higher Order Causality in Complex Systems

### v1

@@ IV Dendritic computation of XOR
We further focus more closely on another example of higher order causal interactions in neuronal dynamics. Despite the long-standing assumption that computing logical XOR requires a circuit of connected neurons, recent work showed that calcium mediated dendritic action potentials (dCaAPs) in a single human cortical neuron can compute this function [20]. In essence, this computation is provided by anti-coincidence behaviour on the apical dendrite of the pyramidal neuron, where two simultaneous synaptic inputs surprisingly reduce dCaAP amplitude (in contrast to the traditional amplification of dendritic AP). We use the biophysical model of dCaAPs provided by [20], simulating its dynamics while dynamically modulating the synaptic inputs coming from two distinct sources, the activity of which corresponded to two parallel sequentially generated logical inputs, where low/high synaptic input represented logical value of 0/1 respectively.
The resulting activity at the apical dendrite (”XOR gate”) is shown in Fig. 1.
@@
To determine causal relationships, the (conditional) mutual information between the input variables ($X_{1},$ $X_{2}$ and $\{X_{1},X_{2}\}$) and the target variable $Y$ was evaluated. As the finite sample estimate of mutual information is generally nonzero, even for independent variables, a statistical test is required. To evaluate the null hypothesis $I(X,Y|Z)=0$ at significance level $\theta=0.01,$ we use a permutation test (using $N=1000$ realizations of shuffled source variable to generate the null distribution).
@@ V Conclusions
We studied the problem of defining the causal effect in systems exhibiting higher-order causal patterns such as XOR-like interactions. We showed that the traditional approaches do not provide sufficiently refined answer. We tackle this challenge by introducing a new, refined definition of causal structure and discuss its relation to the original definition. A trade-off for the increased expressiveness of the hypernetwork definition is its exponential complexity due to formally requiring iteration over a powerset in the condition. This calls for iterative approximation algorithms already commonly utilized for transfer entropy estimation. The main theoretical challenge lies then in treating conveniently the cases of combined synergistic and direct, or even mediating, causal effect.

### v2

@@ IV Dendritic computation of XOR
We further focus more closely on another example of higher order causal interactions in neuronal dynamics. Despite the long-standing assumption that computing logical XOR requires a circuit of connected neurons, recent work showed that calcium mediated dendritic action potentials (dCaAPs) in a single human cortical neuron can compute this function [23]. In essence, this computation is provided by anti-coincidence behaviour on the apical dendrite of the pyramidal neuron, where two simultaneous synaptic inputs surprisingly reduce dCaAP amplitude (in contrast to the traditional amplification of dendritic AP). We use the biophysical model of dCaAPs provided by [23], simulating its dynamics while dynamically modulating the synaptic inputs coming from two distinct sources, the activity of which corresponded to two parallel sequentially generated logical inputs, where low/high synaptic input represented logical value of 0/1 respectively. The resulting activity at the apical dendrite (”XOR gate”) is shown in Fig. 2.
@@
To determine causal relationships, the (conditional) mutual information between the input variables ($X_{1},$ $X_{2}$ and $\{X_{1},X_{2}\}$) and the target variable $Y$ was evaluated. As the finite sample estimate of mutual information is generally nonzero, even for independent variables, a statistical test is required. To evaluate the null hypothesis $I(\mathbf{X}_{I},Y|\mathbf{S})=0$ at significance level $\theta=0.01,$ we use a permutation test (using $N=1000$ realizations of shuffled source variable(s) $\mathbf{X}_{I}$ to generate the null distribution). This breaks any dependency (pairwise or multivariate) between $\mathbf{X}_{I}$ and $Y$ while maintaining dependencies between $Y$ and $\mathbf{S}$.
@@ V Conclusions
We studied the problem of defining the causal effect in systems exhibiting higher-order causal patterns such as XOR-like interactions. We showed that the traditional approaches do not provide sufficiently refined answer. We tackle this challenge by introducing a new, refined definition of causal structure and discuss its relation to the original definition. A trade-off for the increased expressiveness of the causal hypernetwork definition is its exponential complexity due to formally requiring iteration over a powerset in the condition. This calls for iterative approximation algorithms already commonly utilized for transfer entropy estimation. The main theoretical challenge lies then in treating conveniently the cases of combined synergistic and direct, or even mediating, causal effect.